Mr. Zito leaned a 10 foot ladder against the wall, placing the bottom of the ladder on the floor 3 feet away from the wall. How high up on the wall was the ladder?
Answer by mananth(16949)

(hyp)^2-(leg1)^2=(leg2)^2
10^2-3^2= h^2
100-9= =h^2
91=h^2
h=9.54 feet